Circular No. 8561

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 9P/TEMPEL H. U. Keller, S. F. Hviid, and M. Küppers, Max-Planck-Institut für

Sonnensystemforschung; I. Bertini, University of Bern; P. Gutiérrez, Instituto de Astrofísica de Andalucía; L. Jorda, Laboratoire d’Astrophysique de Marseille; J. Knollenberg, Institut für Planetenerkundung, DLR; D. Koschny, European Space Agency (ESA); and the OSIRIS team report that the ‘Optical, Spectroscopic, and Infrared Remote Imaging System’ (OSIRIS) on ESA’s Rosetta spacecraft observed comet 9P/Tempel continuously around the impact of the ‘Deep Impact’ projectile. OSIRIS took images of the dust with the clear and orange (640 ± 43 nm) filters. A sevenfold increase in the brightness of the inner coma (central pixel brightness on a scale of 1500 km/pixel) was measured in the first 30 min after the impact. The brightness of the comet stayed at its enhanced level for at least 8 hr. The expansion velocity of the impact-created dust cloud was 200–250 m/s for the first 8 hr after the impact.

Total visual magnitude estimates: June 9.00 UT, 10.1 (J. J. Gonzalez, Leon, Spain, 25×100 binoculars); 23.89, 10.2 (K. Hornoch, Lelekovice, Czech Republic, 0.13-m reflector); 27.93, 10.3 (A. Baransky, Pylypovychi, Ukraine, 0.36-m reflector); July 1.97, 10.1 (Gonzalez); 3.23, 9.9 (C. E. Spratt, Victoria, BC, 0.08-m refractor; twilight); 5.00, 10.5 (J. G. de S. Aguiar, Campinas, Brazil, 25×100 binoculars); 5.94, 9.9 (Gonzalez).

V5116 SAGITTARII = NOVA SAGITTARII 2005 No. 2 E. Kazarovets and the GCVS team report that the variable-star des-

ignation V5116 Sgr has been assigned to this nova (cf. IAUC 8559).

Additional visual magnitude estimates: July 5.592 UT, 8.2 (A. Pearce, Nedlands, W. Australia); 5.960, 8.6 (P. Schmeer, Bischmisheim, Germany); 5.993, 8.2 (A. Kammerer, Malsch, Germany); 6.476, 8.3 (Pearce); 6.89, 8.4 (M. Lehky, Hradec Kralove, Czech Republic); 6.894, 8.6 (K. Hornoch, Lelekovice, Czech Rep.); 6.896, 8.8 (A. Baransky, Pylypovychi, Ukraine).

COMET C/2005 A1 (LINEAR) Further to IAUC 8559, the June 25 observations of component B were

reported together with the confirming July 4 observations.

2005 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green